OpenStudy (anonymous):

simplify (^x+^3)(^x+^27)

OpenStudy (anonymous):

makes no sense

OpenStudy (anonymous):

whats '^x' ?

OpenStudy (anonymous):

sorry. I guess I need to be more specific. simplify (square root of x+square root of 3)(square root of x+square root of 27)

OpenStudy (jennychan12):

\[(\sqrt{x} + \sqrt{3}) (\sqrt{x}+3\sqrt{3})\] just FOIL

OpenStudy (jennychan12):

@mathhelp7 are you getting it?

OpenStudy (anonymous):

no. pls show me through.

OpenStudy (anonymous):

(a + b ) ( c + d ) = ac + ad + bc + bd

OpenStudy (anonymous):

so is this the answer? (x √ +3 √ )(x √ +3 √3

OpenStudy (anonymous):

no use the formula I posted above and simplify.

OpenStudy (jennychan12):

what @Asad0000 means: a = √x b = √3 c = √x d = 3 √3 plug this stuff into the formula he posted above.

OpenStudy (anonymous):

Oh ok. thank you so much!!!

OpenStudy (anonymous):

can you help me with 2 more problems pls?

OpenStudy (anonymous):

What is the simplest form of the radical expression? √ 2 √ + √5 divide by √3 - √5

OpenStudy (anonymous):

Add 3 divided by 2 +√ 2 + 3 divided by 2 - √5

OpenStudy (jennychan12):

multiply by the conjugate of denominator for all if (a+b). then conjugate is (a-b)

OpenStudy (anonymous):

would you pls show work? I'd appreciate it.

OpenStudy (jennychan12):

do u not get it? cuz if u do, i'd like to see your work (even if you're not sure)

OpenStudy (jennychan12):

√ 2 √ + √5 divide by √3 - √5 ^ is that other √ just a typo?

OpenStudy (anonymous):

sorry. Yes it was a typo. Add 3 divided by 2 +√ 5 + 3 divided by 2 - √5

OpenStudy (jennychan12):



OpenStudy (anonymous):

Im doing two different questions. 1.) What is the simplest form of the radical expression? √ 2 + √5 divide by √2 - √5 2.) Add 3 divided by 2 +√ 5 + 3 divided by 2 - √5

OpenStudy (jennychan12):

okay lemme start u off the first one...  can u try it